We're still chipping away at the data...
Sorry for the delay in responding to your messages.
We have compiled all the data into a single database (no mean feat as it's now 118GB!) due to our limited resources this was spread over a number of HDs. The FM@H data is 95% complete. Kevin is in the process of pulling out the remaining failed dockings for re-docking on the BOINC server.
The new name of the server will be FiND@Home for FightNeglectedDiseases@Home. Over the summer we'll build a website front-end for the data, that will be searchable and interlink with other online resources (e.g. Uniprot/PDB/KEGG). This should make the data easily accessible for biologists interested in exploring the output data. We'll build it with future BOINC data from other genomes in mind (e.g. TB).
Many thanks for hanging in there for us...
ps: We would appreciate it if users do not bad-mouth other projects, as we do not want to have to act as social police. The reason we're NOT concentrating on these boards at the moment is because we're trying to concentrate on getting the data we have into a presentable format for publication. And to focus on validation studies.
14 Mar 2014, 10:56:43 UTC
Errors in current batch or WUs - BACK ONLINE
I'm sure you have noticed errors in the latest batch of WUs. This is due to formatting errors on a few ligands and results in many failed WUs because each ligand is paired with each receptor, therefore 1 failed ligand is muliplied by the number of proteins with which it is docked.
We have always had some errors in our ligands but the rate has been quite low (< 0.5%) overall until now. As the rate has increased (8% in the last 24 hours), I will be disabling the work generator until tomorrow afternoon while I weed out some of the errors.
Thanks for the feedback,
18 Nov 2013, 18:46:47 UTC
We had a power outage over the weekend. The server is back up now and we should be up and running with new WUs in the next 30 mins.
11 Nov 2013, 11:02:54 UTC
No new tasks for 24 hours
Sorry guys but the server is overloaded at the moment. We have stopped submitting new tasks while the daemons catch up. Also, we are going to start a new experiment tomorrow afternoon. Hopefully we will be up and running by 5.
4 Nov 2013, 17:57:50 UTC
Brief interruption to supply of tasks
We are preparing the next experiment now. We expect to have more tasks this evening.
1 Nov 2013, 10:16:01 UTC
News is available as an RSS feed